OpenStudy (anonymous):

Write as a single fraction. -1 +(2c+5d)/(3c) - (c-9d)/(6c)

OpenStudy (anonymous):

119d-3c -------- 6c

OpenStudy (anonymous):

i got d-3c / 6c

zepdrix (zepdrix):

\[\large -1+\frac{(2c+5d)}{3c}-\frac{(c-9d)}{6c}\] Getting a common denominator gives us,\[\large -\frac{6c}{6c}+\frac{2(2c+5d)}{6c}-\frac{(c-9d)}{6c}\] \[\large \frac{-6c+4c+10d-c+9d}{6c}\] \[\large =\frac{19d-3c}{6c}\] @jtabby I think you forgot to distribue the - to each term in the brackets on the last part.\[\large -(c-9d) \quad = \quad -c+9d\]

OpenStudy (anonymous):

ohhh yes i did forget to distribute the negative sign! thank you so much! and how do you make it formatted all nicely like that?

zepdrix (zepdrix):

When you want to enter a message, there are several helpful tools at the bottom. Equation, Draw, Attach File. I entered that text into the Equation tool. It's based off the LaTeX Formatting Language. There are a lot of helpful buttons in the Equation tool, but it takes a little practice to get comfortable with it. :)